OpenStudy (anonymous):

3. Lena Dimock is saving for her college expenses. She sets aside $200 at the beginning of each three months in an account paying 8% annual interest, compounded quarterly. How much will Lena have accumulated in the account at the end of four years?

OpenStudy (saifoo.khan):

$3727.86

OpenStudy (anonymous):

how did you get this answer

OpenStudy (anonymous):

Hold on, I just solve that problem

OpenStudy (saifoo.khan):

And i just answered!

OpenStudy (anonymous):

Here is the derivation of formula: http://openstudy.com/users/imranmeah91#/users/imranmeah91/updates/4e2c32990b8b3d38d3ba31fc

OpenStudy (anonymous):

Here is the formula \[P\left(\frac{\left(1-(1+R)^N\right.}{-R}\right)\]

OpenStudy (anonymous):

that I derived above

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

I got $11270

OpenStudy (anonymous):

i got something different

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

15369

OpenStudy (anonymous):

were you evalutating this :\[200\frac{1-\left(1+\frac{.08}{12}\right)^{48}}{\frac{-.08}{12}}\]

OpenStudy (anonymous):

no did something completely different

OpenStudy (anonymous):

I will check answer using excel

OpenStudy (anonymous):

thank you for your help

OpenStudy (anonymous):

let me know if you have any question, also if you click on the link above , I showed derivation to the equation I used